### v1

# Weak–strong uniqueness for bi-fluid compressible system with algebraic closure

Abstract
We consider a real two-fluid system of compressible viscous fluids with a common velocity field and algebraic closure for the pressure law. The constitutive relation involves densities of both fluids through an implicit function. The existence of global-in-time finite energy weak solutions to this system is known since the work of Novotný and Pokorný [Arch. Rational Mech. Anal., 2020]. On the other hand, existence of local-in-time strong solutions is due to Piasecki and Zatorska [J. Math Fluid Mech., 2022]. In this paper, we establish the weak–strong uniqueness principle using the relative entropy method. In sharp contrast to the two-phase model of Baer-Nunziato type, the volume fraction of phase ++ obeys a transport equation with an additional nonlinear term. This gives rise to troublesome terms in the relative entropy inequality. We are able to close the estimate by making an elaborate use of the structure of the system.
Keywords: Two-fluid model, algebraic closure, global weak solutions, relative entropy.
2020 Mathematics Subject Classification: 35Q35, 76N10.
1 Introduction
Multi-component compressible flow models (often termed multifluid or two-phase models) arise from interface-averaging procedures applied to microscopic balance laws. In the averaged description, volume fractions become additional unknowns and various closure assumptions are needed to close the macroscopic system. Classical modeling expositions may be found in the monographs of Ishii and Hibiki [13] and Drew and Passman [7], as well as in more modern surveys of mathematical literature on multifluid systems [4, 18].
A common modeling starting point is the two-velocity Baer–Nunziato framework [1], from which various one-velocity reductions are obtained by relaxation and closure assumptions [16, 5].
Two analytically distinct one-velocity regimes are:
  • Differential closure: the volume fraction $\alpha$ obeys a transport type equation (typical in one-velocity Baer–Nunziato-type models). This transport equation should include some relaxation term between the pressures, see [4, 5], but simplified models with pure transport equation have also been studied [25, 15].
  • Algebraic closure: the phases share a common velocity and are coupled by an implicit constitutive relation—most prominently equal pressure $p_{+}=p_{-}$. In this setting, $\alpha$ is not an independent transported scalar once one passes to the “transported” partial masses $R=\alpha{\varrho_{+}}$, $Q=(1-\alpha){\varrho_{-}}$; instead, the implicit pressure relation generates additional nonlinear structure, see below for more description.
The present paper analyzes the second class and considers the viscous compressible bi-fluid system in a bounded three-dimensional domain $\Omega$, where both phases are barotropic (power-law pressures) and the single momentum equation is driven by the mixture density $R+Q$.
For mono-fluid compressible Navier–Stokes equations, Lions-Feireisl theory [22, 10] relies crucially on monotonicity/compensated compactness mechanisms tied to a pressure depending on a single conserved quantity. In multifluid systems, the pressure typically depends on several quantities; even when these quantities satisfy the continuity equations, their weak convergences do not in general commute with nonlinear implicit pressure maps. This gives rise to new challenges.
A first major step for algebraic closures in the semi-stationary regime is the global finite-energy existence theory for the two-fluid Stokes system by Bresch, Mucha, and Zatorska [3]. Their analysis makes explicit the core obstruction: after reformulation, the pressure argument satisfies a transport equation with a nonlinear production term, preventing the standard effective-viscous-flux compactness route; this is precisely where the Bresch–Jabin compactness criterion was brought in [2].
The fully time-dependent viscous bi-fluid model with algebraic pressure closure was then treated by Novotný and Pokorný, who proved the existence of global-in-time finite-energy weak solutions for large initial data [26]. Their strategy was to transform the original variables into a two-densities Navier–Stokes-type system with a pressure of “complicated structure dependent on two variable densities”, and then to adapt the Lions–Feireisl method supplemented by DiPerna–Lions transport/renormalization ideas[6] refined in [24, 28].
These results were then extended for differential closures (transport of $\alpha$) by Novotný [25]. In particular, Jin and Novotný [15] established the weak–strong uniqueness property of such solutions, emphasizing the role of convexity of the associated Helmholtz potential. See also [17] for problems with inflow-outflow boundary conditions, and [14] for the analogous stability result but for the dissipative measure-valued solutions. Still, the assumption about convexity of the free Helmholtz energy with respect to the conservative variables seems to be of paramount importance.
On the strong-solution side, local-in-time well-posedness for the compressible two-fluid algebraic-closure system was obtained by Piasecki and Zatorska using the maximal-regularity techniques [27]. The existence of global-in-time strong solutions for small perturbations of constant initial conditions is due to Evje, Wang and Wen [8].
Given the coexistence of global weak solutions and local strong solutions, a natural stability question is whether weak solutions are unique as long as a strong solution exists—the weak–strong uniqueness principle. For the mono-fluid compressible Navier–Stokes system, weak–strong uniqueness via the relative entropy method is nowadays classical [11], and has proven to have profound applications in the study of singular-limits problems, see for example [12, 23]. For the two-fluid system, the only known results, as far as we are aware, are aforementioned [15, 14] and a conditional result by Li and Zatorska [21]. The singular-limit problems, including for example the low Mach number limit, are therefore still mostly formal [18], however some recent progress for multi-component flows of MHD type can be found in [9].
In sharp contrast to differential-closure models, the algebraic constraint ${p_{+}}={p_{-}}$ changes the evolution of the volume fraction to transport with an additional nonlinear term. This phenomenon appears already in the Stokes analysis of Bresch–Mucha–Zatorska [3] and is precisely what breaks the most direct relative-entropy estimates: extra terms arise that do not cancel as in the pure-transport case. The goal of this paper is to prove the weak–strong uniqueness principle for the viscous compressible bi-fluid system with algebraic pressure closure: any finite-energy weak solution in the sense of Novotný–Pokorný [26] coincides, on the lifespan of the strong solution, with the local strong solution constructed by Piasecki–Zatorska [27] emanating from the same initial data. Our proof follows the relative entropy strategy but requires an additional mechanism to control the defect in the volume fraction generated by the algebraic closure. More concretely, we exploit the specific implicit structure of the pressure law to derive a stability estimate for solution compatible with the relative energy inequality and then conclude by a Grönwall argument. It should be emphasized that, in sharp contrast with Jin–Novotný [15], no positive bound for $\alpha$ from below and away from $1$ is required in our analysis.
The paper is organized as follows. In Section 2 after recalling the model and the notion of bounded-energy weak solutions, we summarize the known existence theory for global weak solutions and local strong solutions. We then derive the relative energy inequality, highlighting the additional nonlinear terms caused by the evolution law for the volume fraction in Section 3. Next, in Section 4 we establish the relative entropy inequality with the strong solution as a reference state and estimate the resulting remainder terms using the structure induced by the algebraic closure. Finally, in Section 5, we combine these estimates to prove the weak–strong uniqueness principle, which is the main achievement of our paper.
2 Formulation of the problem
Our system of equations reads:
$\partial_{t}(\alpha{\varrho_{+}})+\operatorname{div}(\alpha{\varrho_{+}}{\bf u})=0,$ (2.1a)
$\partial_{t}((1-\alpha){\varrho_{-}})+\operatorname{div}((1-\alpha){\varrho_{-}}{\bf u})=0,$ (2.1b)
$\partial_{t}((\alpha{\varrho_{+}}+(1-\alpha){\varrho_{-}}){\bf u})+\operatorname{div}((\alpha{\varrho_{+}}+(1-\alpha){\varrho_{-}}){\bf u}\otimes{\bf u})$
$\qquad\qquad\qquad\qquad\qquad+\nabla(\alpha p_{+}({\varrho_{+}})+(1-\alpha)p_{-}({\varrho_{-}}))=\operatorname{div}{\bf S}(\nabla{\bf u}),$ (2.1c)
$p_{+}({\varrho_{+}})=p_{-}({\varrho_{-}}).$ (2.1d)
In the system above, we denoted ${\bf S}={\bf S}(\nabla{\bf u})$ to be the usual Newtonian stress tensor given by
$\operatorname{div}{\bf S}(\nabla{\bf u})=\mu\Delta{\bf u}+(\mu+\lambda)\nabla\operatorname{div}{\bf u},$
where the two constants $\mu,\lambda$ are viscosity coefficients. We take $\mu>0$ and $2\mu+3\lambda\geq 0$. By $_{+}, pp_{-}$ we denote the internal barotropic pressures for each fluid with the explicit form:
$\begin{split}p_{+}=({\varrho_{+}})^{\gamma^{+}},\quad p_{-}=({\varrho_{-}})^{\gamma^{-}},\end{split}$ (2.2)
where $\gamma^{\pm}>1$ are given adiabatic exponents. The system is endowed with boundary condition
${\bf u}(t,x)\Big|_{\partial\Omega}={\bf 0}$ (2.3)
on $(0,T)\times\partial\Omega$, where $\Omega$ is a sufficiently smooth bounded domain of $\mathbb{R}^{3}$ and $T>0$ is arbitrary large. We consider initial conditions in $\Omega$
$\alpha\varrho_{+}(0,x)$ $=\alpha_{0}\varrho_{+,0}(x)=:R_{0}(x),$ (2.4)
$(1-\alpha)\varrho_{-}(0,x)$ $=(1-\alpha_{0})\varrho_{-,0}(x)=:Q_{0}(x),$
$(\alpha\varrho_{+}+(1-\alpha)\varrho_{-}){\bf u}(0,x)$ $=(\alpha_{0}\varrho_{+,0}+(1-\alpha_{0})\varrho_{-,0}){\bf u}_{0}(x)=:{\bf m}_{0}.$
The main result of this paper reads
Theorem 2.1
Let $(\alpha,\varrho_{-},\varrho_{+},{\bf u})$ be a global weak solution to the system (2.1)–(2.4) obtained in [26], and let $(\beta,{\widetilde{\varrho}_{-}},{\widetilde{\varrho}_{+}},{\bf v})$ be the strong solution obtained in [27] emanating from the same initial data. Then the two solutions coincide on the time interval $(0,T)$ of a lifespan of the regular solution.
To prove this theorem, we use the method of relative entropy, inspired by Jin and Novotný [15]. We will show that the following relative entropy functional
$\begin{split}&{\cal E}(\alpha,R,Q,{\bf u}\,|\,\beta,{\widetilde{R}},{\widetilde{Q}},{\bf v})(t)\\ =&\int_{\Omega}\frac{1}{2}(R+Q)|{\bf u}-{\bf v}|^{2}(t,x){\rm d}{x}+\int_{\Omega}\frac{1}{2}(\alpha-\beta)^{2}(t,x){\rm d}{x}\\ &+\int_{\Omega}\alpha\Big({H_{+}}({\varrho_{+}})-{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\varrho_{+}}-{\widetilde{\varrho}_{+}})-{H_{+}}({\widetilde{\varrho}_{+}})\Big)(t,x){\rm d}{x}\\ &+\int_{\Omega}(1-\alpha)\Big({H_{-}}({\varrho_{-}})-{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})({\varrho_{-}}-{\widetilde{\varrho}_{-}})-{H_{-}}({\widetilde{\varrho}_{-}})\Big)(t,x){\rm d}{x},\end{split}$ (2.5)
where for brevity we set
[…]
and
$H_{+}({\varrho_{+}})=\frac{1}{{\gamma^{+}}-1}{\varrho_{+}}^{\gamma^{+}},\quad{H_{-}}({\varrho_{-}})=\frac{1}{{\gamma^{-}}-1}{\varrho_{-}}^{\gamma^{-}},$
remains small, provided it is small initially, and that the regular solution $(\beta,{\widetilde{R}},{\widetilde{Q}},{\bf v})$ emanating from the same initial data exists.
2.1 Global weak solutions
We now recall the definition of a weak solution to the system (2.1), see Definition 2 from Novotný and Pokorný [26], and the corresponding existence Theorem 2. For notational convenience, we set $I:=(0,T)$.
Definition 1
The quadruple $(\alpha,\varrho_{-},\varrho_{+},{\bf u})$ is called a bounded energy weak solution to the problem (2.1)–(2.4) in $I\times\Omega$, if $0\leq\alpha\leq 1$, $\varrho_{\pm}\geq 0$ $a.$ in $I\times\Omega$, $\varrho_{\pm}\in L^{\infty}(I;L^{1}(\Omega))$, ${\bf u}\in L^{2}(I;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}))$, $(\alpha\varrho_{+}+(1-\alpha)\varrho_{-})|{\bf u}|^{2}\in L^{\infty}(I;L^{1}(\Omega))$, ${p_{-}}({\varrho_{-}})={p_{+}}({\varrho_{+}})\in L^{1}(I\times\Omega)$, and:
  • Continuity equations (2.1a) and (2.1b) are satisfied in the weak sense
    $\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(\alpha{\varrho_{+}}\partial_{t}\psi+\alpha{\varrho_{+}}{\bf u}\cdot\nabla\psi\big)\ {\rm d}{x}\,{\rm d}t+\int_{\Omega}R_{0}\psi(0,\cdot)\ {\rm d}{x}=0,$ (2.7)
    $\int_{0}^{T}\!\!\!\!\int_{\Omega}\big((1-\alpha){\varrho_{-}}\partial_{t}\psi+(1-\alpha){\varrho_{-}}{\bf u}\cdot\nabla\psi\big)\ {\rm d}{x}\,{\rm d}t+\int_{\Omega}Q_{0}\psi(0,\cdot){\rm d}{x}=0,$ (2.8)
    for any $\psi\in C^{1}_{c}([0,T)\times\overline{\Omega})$.
  • Momentum equation (2.1c) is satisfied in the weak sense
    $\int_{0}^{T}\!\!\!\!\int_{\Omega}\big((\alpha{\varrho_{+}}+(1-\alpha){\varrho_{-}}){\bf u}\cdot\partial_{t}{\boldsymbol{\varphi}}+(\alpha{\varrho_{+}}+(1-\alpha){\varrho_{-}})({\bf u}\otimes{\bf u}):\nabla{\boldsymbol{\varphi}}+p\operatorname{div}{\boldsymbol{\varphi}}\big)\ {\rm d}{x}\,{\rm d}t$ (2.9)
    $=$ $\int_{0}^{T}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}):\nabla{\boldsymbol{\varphi}}\ {\rm d}{x}\,{\rm d}t-\int_{\Omega}{\bf m}_{0}\cdot{\boldsymbol{\varphi}}(0,\cdot)\ {\rm d}{x},$
    where $p=\alpha p_{+}({\varrho_{+}})+(1-\alpha)p_{-}({\varrho_{-}})$, and for any ${\boldsymbol{\varphi}}\in C^{1}_{c}([0,T)\times\Omega;\mathbb{R}^{3})$.
  • The energy estimate
    $\begin{split}&\int_{\Omega}\left(\frac{1}{2}(\alpha{\varrho_{+}}+(1-\alpha){\varrho_{-}})|{\bf u}|^{2}+\alpha{H_{+}}({\varrho_{+}})+(1-\alpha){H_{-}}({\varrho_{-}})\right)(\tau)\ {\rm d}{x}\\ &\qquad\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}):\nabla{\bf u}\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{\Omega}\left(\frac{1}{2}(\alpha{\varrho_{+}}+(1-\alpha){\varrho_{-}})|{\bf u}|^{2}+\alpha{H_{+}}({\varrho_{+}})+(1-\alpha){H_{-}}({\varrho_{-}})\right)(0)\ {\rm d}{x}\end{split}$ (2.10)
    holds for a.a. $\tau\in I$.
Under the algebraic closure for the pressure, it has been proven by Novotný and Pokorný [26] (see Section 5 therein) that the system (2.1) admits global-in-time weak solutions.
Theorem 2.2
Let $0\leq\underline{a}<\overline{a}<\infty$, $\gamma_{BOG}^{\pm}=\min\{\frac{2}{3}\gamma^{\pm}-1,\frac{\gamma^{\pm}}{2}\}$. Let $G:=\gamma^{+}+\gamma^{+}_{BOG}$ if $\underline{a}=0$ and $G:=\max\{\gamma^{+}+\gamma^{+}_{BOG},\gamma^{-}+\gamma^{-}_{BOG}\}$ if $\underline{a}>0$. Assume
$0<\gamma^{-}<\infty,\qquad\gamma^{+}\geq\frac{9}{5},\qquad\overline{\Gamma}<G,$ (2.11)
where
$\overline{\Gamma}:=\left\{\begin{array}[]{c}\max\{\gamma^{+}-\frac{\gamma^{+}}{\gamma^{-}}+1,\,\gamma^{-}+\frac{\gamma^{-}}{\gamma^{+}}-\frac{\gamma^{+}}{\gamma^{-}}\}\qquad\mbox{if $\underline{a}=0$}\\ \max\{\gamma^{+}-\frac{\gamma^{+}}{\gamma^{-}}+1,\gamma^{-}+\frac{\gamma^{-}}{\gamma^{+}}-1\}\qquad\mbox{if $\underline{a}>0$}\end{array}\right\}.$
Suppose that
$0\leq\alpha_{0}\leq 1,\qquad\underline{a}\alpha_{0}\varrho_{+,0}\leq(1-\alpha_{0})\varrho_{-,0}\leq\overline{a}\alpha_{0}\varrho_{+,0},\qquad p_{+}(\varrho_{+,0})=p_{-}(\varrho_{-,0}),$ (2.12)
$\varrho_{+,0}\in L^{\gamma^{+}}(\Omega),\qquad{\frac{|{\bf m}_{0}|^{2}}{\alpha_{0}\varrho_{+,0}+(1-\alpha_{0})\varrho_{-,0}}\in L^{1}(\Omega).}$
Then the problem (2.1)–(2.4) admits at least one weak solution in the sense of Definition 1. Moreover, $\alpha\varrho_{+}$ belongs to the space $C_{\rm{weak}}([0,T];L^{\gamma^{+}}(\Omega))$ and $(1-\alpha)\varrho_{-}$ belongs to the space $C_{\rm{weak}}([0,T];L^{q_{\gamma^{+},\gamma^{-}}}(\Omega))$ [^1] Here, the notation is $q_{\theta_{1},\theta_{2}}=\theta_{1}$ if $\theta_{2}<\theta_{1}$ and $q_{\theta_{1},\theta_{2}}=\theta_{2}$ if $\theta_{2}\geq\theta_{1}$., the vector field $(\alpha\varrho_{+}+(1-\alpha)\varrho_{-}){\bf u}$ belongs to $C_{\rm{weak}}([0,T];L^{r}(\Omega;\mathbb{R}^{3}))$ for some $r>1$, and $p_{\pm}(\varrho_{\pm})\in L^{r}((0,T)\times\Omega)$ for some $r>1$.
2.2 Local-in-time strong solutions
Here we recall the local well-posedness for our bi-fluid system with large initial data from the work of Piasecki and Zatorska[27]. To this purpose, we consider the following reformulation of the system (2.1) with $R$ and $Q$ defined in (2.6)
$\partial_{t}R+\operatorname{div}(R{\bf u})=0,$ (2.13a)
$\partial_{t}Q+\operatorname{div}(Q{\bf u})=0,$ (2.13b)
$\partial_{t}((R+Q){\bf u})+\operatorname{div}((R+Q){\bf u}\otimes{\bf u})-\operatorname{div}{\bf S}({\bf u})+\nabla Z^{\gamma^{+}}={\bf 0}.$ (2.13c)
In the system above the pressure $p$ is expressed in terms of $R,Q$. In fact we have
$p=P(R,Q)={Z}^{\gamma_{+}},$ (2.14)
for $Z=Z(R,Q)$ such that
$\begin{split}Q=\left(1-\frac{R}{Z}\right)Z^{\gamma},\quad\mbox{with}\quad\gamma=\frac{\gamma_{+}}{\gamma_{-}},\end{split}$ (2.15)
and
$\begin{split}R\leq Z.\end{split}$ (2.16)
Relation (2.14) identifies $Z$ as $\varrho_{+}$} and thus, (2.16) can be used to deduce existence of $\alpha$ such that $0\leq\alpha\leq 1$ and $R=\alpha Z$, see [3] and also [26] for a similar argument.
The existence of local-in-time strong solutions to the system (2.13) is stated below.
Theorem 2.3
Assume that $\Omega$ is a uniform $C^{2}$ domain and $2<p<\infty,\;3<q<\infty,\;\frac{2}{p}+\frac{3}{q}<1$. Assume moreover that $R_{0},Q_{0}$ satisfy
${R}_{0}\geq 0,\quad{Q}_{0}\geq 0,\quad R_{0}+Q_{0}\geq\kappa\;\;\textrm{for some}\;\kappa>0$
and ${\bf u}_{0}$ satisfies the compatibility condition
${\bf u}_{0}|_{\partial\Omega}=\mathbf{0}.$
Then for any $L>0$ there exists $T>0$ such that if
$\|\nabla R_{0}\|_{L^{q}(\Omega)}+\|\nabla Q_{0}\|_{L^{q}(\Omega)}+\|{\bf u}_{0}\|_{B^{2-2/p}_{q,p}(\Omega)}\leq L,$
then (2.13) supplemented by the boundary condition (2.3), the initial conditions (2.4) and the relation (2.15) admits a unique solution $(R,Q,{\bf u})$ on $(0,T)$ with the estimate
$\|(R,Q,{\bf u})\|_{{\mathcal{X}}(T)}\leq CL,\qquad\int_{0}^{T}\|\nabla{\bf u}(t)\|_{L^{\infty}(\Omega)}\,{\rm d}t\leq\delta,$
where
$\|(g,h,{{\bf F}})\|_{{\cal{X}}(T)}:=\|{{\bf F}}\|_{L^{p}(0,T;W^{2,q}(\Omega))}+\|{{\bf F}}_{t}\|_{L^{p}(0,T;L^{q}(\Omega))}+\|(g,h)\|_{W^{1,p}(0,T;W^{1,q}(\Omega))}$
and $\delta$ is a small positive constant.
Remark 2.4
Although Theorem 2.3 provides a local well-posedness for the reformulation of the system (2.1) in which $R=\alpha{\varrho_{+}}$, $Q=(1-\alpha){\varrho_{-}}$ and $p=p(R,Q)$, due to the algebraic closure (2.1d), $\alpha$ can be recovered as $\alpha(R,Q)$, see the discussions in Bresch–Mucha–Zatorska [3]. Hence, the original system (2.1) admits a unique local strong solution for large initial data.
Remark 2.5
Theorem 2.3 provides existence of solutions in the regularity class required for the purposes of the relative entropy argument. In particular, one can verify that
$\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{1}(0,T;L^{\infty}(\Omega))}+\|{\bf v}\|_{L^{1}(0,T;W^{1,\infty}(\Omega))}\leq C,$
$\|{\varrho_{+}}\|_{L^{\infty}(0,T;W^{1,\infty}(\Omega))}+\|{\varrho_{-}}\|_{L^{\infty}(0,T;W^{1,\infty}(\Omega))}\leq C,$
as well as $\alpha$ such that
$\|\alpha\|_{L^{\infty}(0,T;W^{1,\infty}(\Omega))}\leq C.$
3 Relative entropy inequality–preliminaries
In the following parts of this section, we will aim at writing an equation for evolution of the relative energy for $\beta,{\widetilde{R}},{\widetilde{Q}}$ and $\bf v$} sufficiently regular. The distinct difference with respect to the two-fluid model of Baer-Nunziato type [15] is that in our case, the equations for $\varrho_{+}$ and $\varrho_{-}$ are not just the continuity equations, but are more complicated. This follows from the fact that $\alpha$ does not satisfy a pure transport equation, see (3.4) for more details.
3.1 Derivation of the equations for $\alpha$, $\varrho_{+}$ and $\varrho_{-}$}
To compute this relative entropy inequality, we will need the equations for $\alpha$, $\varrho_{+}$} and $\varrho_{-}$}.
Formally, we can calculate
$\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha$ $=\partial_{R}\alpha\partial_{t}R+\partial_{Q}\alpha\partial_{t}Q+\partial_{R}\alpha{\bf u}\cdot\nabla R+\partial_{Q}\alpha{\bf u}\cdot\nabla Q$
$=\partial_{R}\alpha(\partial_{t}R+{\bf u}\cdot\nabla R)+\partial_{Q}\alpha(\partial_{t}Q+{\bf u}\cdot\nabla Q)$
$=-\partial_{R}\alpha R\operatorname{div}{\bf u}-\partial_{Q}\alpha Q\operatorname{div}{\bf u};$
whence
$\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha+(\partial_{R}\alpha R+\partial_{Q}\alpha Q)\operatorname{div}{\bf u}=0.$
Denoting
$\begin{split}\gamma=\gamma^{+}/\gamma^{-}>0\end{split}$ (3.1)
and using the equality of pressures (2.1d) and the constitutive relations (2.2), we can find that
$\begin{split}R^{\gamma}(1-\alpha)=Q\alpha^{\gamma}.\end{split}$ (3.2)
Hence, applying $\partial_{Q}$ and $\partial_{R}$ on both sides, we obtain
$\begin{split}\partial_{Q}\alpha=\frac{-\alpha^{\gamma}}{Q\gamma\alpha^{\gamma-1}+R^{\gamma}},\qquad\partial_{R}\alpha=\frac{\gamma R^{\gamma-1}(1-\alpha)}{Q\gamma\alpha^{\gamma-1}+R^{\gamma}}.\end{split}$ (3.3)
Based on (3.2)-(3.3), it holds
$\partial_{R}\alpha R+\partial_{Q}\alpha Q$ $=\frac{\gamma R^{\gamma}(1-\alpha)}{Q\gamma\alpha^{\gamma-1}+R^{\gamma}}+\frac{-Q\alpha^{\gamma}}{Q\gamma\alpha^{\gamma-1}+R^{\gamma}}$
$=\frac{\gamma R^{\gamma}(1-\alpha)-Q\alpha^{\gamma}}{Q\gamma\alpha^{\gamma-1}+R^{\gamma}}$
$=\frac{(\gamma-1)Q\alpha^{\gamma}}{\frac{\gamma}{\alpha}Q\alpha^{\gamma}+\frac{1}{1-\alpha}Q\alpha^{\gamma}}$
$=\frac{\gamma-1}{\frac{\gamma}{\alpha}+\frac{1}{1-\alpha}}$
$=\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha(1-\alpha).$
Thus we obtain
$\begin{split}\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha+\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha(1-\alpha)\operatorname{div}{\bf u}=0.\end{split}$ (3.4)
Equation (3.4) is of course not exactly a pure transport equation, but we should notice that because $0\leq\alpha\leq 1$, the coefficient
$\omega_{\alpha}:=\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha(1-\alpha)$
is also bounded in $L^{\infty}((0,T)\times\Omega)$. Indeed, $\gamma(1-\alpha)+\alpha$ has a strictly positive lower bound, namely $\gamma(1-\alpha)+\alpha\geq\min\{1,\gamma\}>0$. Thus $|\omega_{\alpha}|\leq\frac{\gamma+1}{\min\{1,\gamma\}}$.
In a similar manner, we obtain
$\begin{split}\partial_{t}(1-\alpha)+{\bf u}\cdot\nabla(1-\alpha)-\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha(1-\alpha)\operatorname{div}{\bf u}=0.\end{split}$ (3.5)
From the equation (3.4), and (2.1a), (2.1b) we now deduce
$\begin{split}\alpha\partial_{t}{\varrho_{+}}+{\varrho_{+}}\partial_{t}\alpha+\operatorname{div}({\varrho_{+}}{\bf u})\alpha+{\varrho_{+}}({\bf u}\cdot\nabla\alpha)=0,\end{split}$ (3.6)
therefore, using (3.4), we obtain
$\partial_{t}{\varrho_{+}}+\operatorname{div}({\varrho_{+}}{\bf u})={\varrho_{+}}\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}(1-\alpha)\operatorname{div}{\bf u}$
and
$\partial_{t}{\varrho_{-}}+\operatorname{div}({\varrho_{-}}{\bf u})=-{\varrho_{-}}\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha\operatorname{div}{\bf u}.$
In order to derive the equations (3.4) and (3.5) rigorously, it suffices to follow the arguments from Section 8.1. in [24], see also Lemma 2.4 in [3]. In particular, we have the following result.
Lemma 3.1
Let ${\bf u}\in L^{2}(0,T;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}))$, $R=\alpha{\varrho_{+}}\in L^{\infty}(0,T;L^{2}(\Omega))$, and $Q=(1-\alpha){\varrho_{-}}\in L^{\infty}(0,T;L^{2}(\Omega))$, and let $(R,Q,{\bf u})$ solve (2.13a) and (2.13b) in the sense of distributions. Then $Z$ defined by (2.15) belongs to $L^{\infty}(0,T;L^{2}(\Omega))$ and it satisfies
$\begin{split}\partial_{t}Z+\operatorname{div}(Z{\bf u})+\frac{(1-\gamma)(Z-R)Z}{\gamma(Z-R)+R}\operatorname{div}{\bf u}=0,\end{split}$ (3.7)
in the sense of distributions.
Conversely, let $(R,Z,{\bf u})$ solve (2.13a) and (3.7) in the sense of distributions. Then $Q$ defined by (2.15) satisfies (2.13b) in the sense of distributions.
The proof of Lemma 3.1 can be found in [3], see Lemma 2.4 and is based on the DiPerna-Lions renormalization theory [6].
Remark 3.2
Let us remark that the assumptions of Lemma 3.1 are satisfied. Indeed, due to the assumption (2.11) and restriction of the initial data (2.12), one can show that the solutions to the continuity equations for $R$ and $Q$ remain in $L^{\infty}(0,T;L^{2}(\Omega))$.
We can now use Lemma 3.1 to identify the equations satisfied by $\alpha$, $\varrho_{+}$} and $\varrho_{-}$}.
Lemma 3.3
Let assumptions of Lemma 3.1 be satisfied. Then $\varrho_{+}$} belongs to $L^{\infty}(0,T;L^{2}(\Omega))$ and it satisfies
$\begin{split}\partial_{t}{\varrho_{+}}+\operatorname{div}({\varrho_{+}}{\bf u})={\varrho_{+}}\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}(1-\alpha)\operatorname{div}{\bf u}\end{split}$ (3.8)
in the sense of distributions. Moreover, $\alpha$ belongs to $L^{\infty}((0,T)\times\Omega)$ and satisfies
$\begin{split}\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha+\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha(1-\alpha)\operatorname{div}{\bf u}=0\end{split}$ (3.9)
in the sense of distributions.
Proof. Let us notice that equation (3.8) follows directly from (3.7) after substituting $Z={\varrho_{+}}$ and $R=\alpha{\varrho_{+}}$. Equation for $\alpha$ is derived the same way thanks to the decomposition (3.6). The bound $0\leq\alpha\leq 1$ follows from the existence result, see Theorem $2.$ $\Box$
3.2 Derivation of relative entropy inequality
We now derive the estimate of a part of the relative entropy functional from (LABEL:rel_entr), namely
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})(\tau)\\ &=\int_{\Omega}\frac{1}{2}(R+Q)|{\bf u}-{\bf v}|^{2}(\tau,x){\rm d}{x}\\ &\quad+\int_{\Omega}\alpha\Big({H_{+}}({\varrho_{+}})-{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\varrho_{+}}-{\widetilde{\varrho}_{+}})-{H_{+}}({\widetilde{\varrho}_{+}})\Big)(\tau,x){\rm d}{x}\\ &\quad+\int_{\Omega}(1-\alpha)\Big({H_{-}}({\varrho_{-}})-{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})({\varrho_{-}}-{\widetilde{\varrho}_{-}})-{H_{-}}({\widetilde{\varrho}_{-}})\Big)(\tau,x){\rm d}{x},\end{split}$ (3.10)
where $(\beta,{\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}},{\bf v})$ are smooth test functions with ${\bf v}|_{\partial\Omega}=\mathbf{0}$, and we set ${\widetilde{R}}=\beta{\widetilde{\varrho}_{+}},\,\,{\widetilde{Q}}=(1-\beta){\widetilde{\varrho}_{-}}$.
We now expand
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}\\ &=\int_{\Omega}\left(\frac{1}{2}(R+Q)|{\bf u}|^{2}+\alpha{H_{+}}({\varrho_{+}})+(1-\alpha){H_{-}}({\varrho_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}\\ &\quad+\int_{\Omega}\frac{1}{2}(R+Q)|{\bf v}|^{2}\ {\rm d}{x}\Big|_{0}^{\tau}\\ &\quad-\int_{\Omega}(R+Q){\bf u}\cdot{\bf v}\ {\rm d}{x}\Big|_{0}^{\tau}\\ &\quad-\int_{\Omega}\alpha{\varrho_{+}}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\Big|_{0}^{\tau}+\int_{\Omega}\alpha\left({H_{+}}^{\prime}({\widetilde{\varrho}_{+}}){\widetilde{\varrho}_{+}}-{H_{+}}({\widetilde{\varrho}_{+}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}\\ &\quad-\int_{\Omega}(1-\alpha){\varrho_{-}}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\Big|_{0}^{\tau}+\int_{\Omega}(1-\alpha)\left({H_{-}}^{\prime}({\widetilde{\varrho}_{-}}){\widetilde{\varrho}_{-}}-{H_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}\end{split}$ (3.11)
and we will treat the right-hand side of (LABEL:rel_form) line by line.
Step 1. The first line is estimated using the energy estimate (LABEL:energy), from which it follows that
$\begin{split}&\int_{\Omega}\left(\frac{1}{2}(R+Q)|{\bf u}|^{2}+\alpha{H_{+}}({\varrho_{+}})+(1-\alpha){H_{-}}({\varrho_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}\leq-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}):\nabla{\bf u}\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.12)
Step 2. We sum the equations for $Q$ and $R$ and test the obtained result by $\frac{1}{2}|{\bf v}|^{2}$
$\begin{split}\int_{\Omega}\frac{1}{2}(R+Q)|{\bf v}|^{2}\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(R+Q){\bf v}\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.13)
Step 3. We take $\bf v$} as a test function in the momentum equation (2.9), and integrate by parts to get
$\begin{split}-\int_{\Omega}(R+Q){\bf u}\cdot{\bf v}\ {\rm d}{x}\Big|_{0}^{\tau}=&-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left((R+Q){\bf u}\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)\right)\ {\rm d}{x}\,{\rm d}t\\ &\qquad\qquad-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(p\operatorname{div}{\bf v}-{\bf S}(\nabla{\bf u}):\nabla{\bf v}\right)\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.14)
[…]
$\begin{split}-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\Big|_{0}^{\tau}=-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}\left(\partial_{t}+{\bf u}\cdot\nabla\right){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\end{split}$ (3.15)
$\begin{split}-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\Big|_{0}^{\tau}=-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}\left(\partial_{t}+{\bf u}\cdot\nabla\right){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.16)
[…]
$\int_{\Omega}\alpha\left({H_{+}}^{\prime}({\widetilde{\varrho}_{+}}){\widetilde{\varrho}_{+}}-{H_{+}}({\widetilde{\varrho}_{+}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}(\alpha{p_{+}}({\widetilde{\varrho}_{+}}))\ {\rm d}{x}\,{\rm d}t.$
In a similar manner
$\int_{\Omega}(1-\alpha)\left({H_{-}}^{\prime}({\widetilde{\varrho}_{-}}){\widetilde{\varrho}_{-}}-{H_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\Big|_{0}^{\tau}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}((1-\alpha){p_{-}}({\widetilde{\varrho}_{-}}))\ {\rm d}{x}\,{\rm d}t.$
Step 6. Summing the above expressions we obtain all together:
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)\ {\rm d}{x}\,{\rm d}t-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}p\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}(\partial_{t}+{\bf u}\cdot\nabla){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}(\partial_{t}+{\bf u}\cdot\nabla){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.17)
Adding to both sides $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t$, we get
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)+{\bf S}(\nabla{\bf v}):\nabla({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}p\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}(\partial_{t}+{\bf u}\cdot\nabla){H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}(\partial_{t}+{\bf u}\cdot\nabla){H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\,{\rm d}t.\end{split}$ (3.18)
[…]
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}}))\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\quad=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\nabla\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\widetilde{\varrho}_{+}}\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\widetilde{\varrho}_{-}}\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t,$
and similarly
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\left(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\right)\ {\rm d}{x}\,{\rm d}t$
$\quad=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\partial_{t}\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\partial_{t}(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\widetilde{\varrho}_{+}}\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\widetilde{\varrho}_{-}}\partial_{t}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t.$
Substituting for these terms in (LABEL:sumstep1b), we obtain
$\begin{split}&\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t\\ &\leq\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big((R+Q)({\bf v}-{\bf u})\cdot\left(\partial_{t}{\bf v}+({\bf u}\cdot\nabla){\bf v}\right)+{\bf S}(\nabla{\bf v}):\nabla({\bf v}-{\bf u})\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}})\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}})\partial_{t}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\alpha{p_{+}}({\widetilde{\varrho}_{+}})+(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})-p\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}{\bf v}-{\varrho_{+}}{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}{\bf v}-{\varrho_{-}}{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\nabla\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\Big)\cdot{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}\alpha{p_{+}}({\widetilde{\varrho}_{+}})+\partial_{t}(1-\alpha){p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t\\ &=:I+II+III+\ldots+VII+VIII.\end{split}$ (3.19)
This is an analogous form of the relative entropy inequality as the one derived in Theorem 5.2 from [15]. Let us stress that up until now we did not use any properties of $(\beta,{\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}},{\bf v})$ other than the sufficient regularity in order to use them as the test functions.
4 Relative entropy inequality for the strong solution
We will now restrict our class of test functions to $(\beta,{\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}},{\bf v})$ given in Theorem 2.3. We reformulate and reduce each of the terms from (LABEL:sumstep1c). In particular,
$I=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left[(R+Q-{\widetilde{R}}-{\widetilde{Q}})\partial_{t}{\bf v}+\left((R+Q){\bf u}-({\widetilde{R}}+{\widetilde{Q}}){\bf v}\right)\cdot\nabla{\bf v}\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t$
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla(\beta{p_{+}}({\widetilde{\varrho}_{+}})+(1-\beta){p_{-}}({\widetilde{\varrho}_{-}}))\ {\rm d}{x}\,{\rm d}t$
$=:$ $\,I_{1}+I_{2}.$
Moreover, we can write
$\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})={H_{+}}^{\prime\prime}({\widetilde{\varrho}_{+}})\partial_{t}{\widetilde{\varrho}_{+}}$
$=-{H_{+}}^{\prime\prime}({\widetilde{\varrho}_{+}})\operatorname{div}({\widetilde{\varrho}_{+}}{\bf v})+{\widetilde{\varrho}_{+}}{H_{+}}^{\prime\prime}({\widetilde{\varrho}_{+}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}(1-\beta)\operatorname{div}{\bf v}$
$=-\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\cdot{\bf v}-{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\operatorname{div}{\bf v}+{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}(1-\beta)\operatorname{div}{\bf v},$
and similarly
$\partial_{t}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})={H_{-}}^{\prime\prime}({\widetilde{\varrho}_{-}})\partial_{t}{\widetilde{\varrho}_{-}}$
$=-{H_{-}}^{\prime\prime}({\widetilde{\varrho}_{-}})\operatorname{div}({\widetilde{\varrho}_{-}}{\bf v})-{\widetilde{\varrho}_{-}}{H_{-}}^{\prime\prime}({\widetilde{\varrho}_{-}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}$
$=-\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\cdot{\bf v}-{p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\operatorname{div}{\bf v}-{p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}.$
Therefore, the terms $II$ and $III$ can be rewritten as
$\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}})\partial_{t}{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$\quad=-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){\bf v}\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$\qquad-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}(1-\beta)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\quad=:II_{1}+II_{2}+II_{3}$
and
$\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}})\partial_{t}{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$\quad=-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){\bf v}\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$\qquad-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\qquad-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\quad=:III_{1}+III_{2}+III_{3}.$
This means we have
$IV+II_{2}+III_{2}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\Big({p_{+}}({\widetilde{\varrho}_{+}})-{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\widetilde{\varrho}_{+}}-{\varrho_{+}})-{p_{+}}({\varrho_{+}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\qquad\qquad\qquad\qquad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)\Big({p_{-}}({\widetilde{\varrho}_{-}})-{p_{-}}^{\prime}({\widetilde{\varrho}_{-}})({\widetilde{\varrho}_{-}}-{\varrho_{-}})-{p_{-}}({\varrho_{-}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t,$
$V+II_{1}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{\varrho_{+}}({\bf v}-{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t,$
$VI+III_{1}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha){\varrho_{-}}({\bf v}-{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t.$
We can now simplify further by writing
$V+II_{1}+VI+$ $III_{1}+I_{2}$
$=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})(\alpha{\varrho_{+}}-\beta{\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$\quad+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\Big((1-\alpha){\varrho_{-}}-(1-\beta){\widetilde{\varrho}_{-}}\Big)\ {\rm d}{x}\,{\rm d}t$
$\quad-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta){p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t.$
Finally, for the terms $VII$ and $VIII$ we keep them at the present form.
[…]
$\begin{split}&{\cal R}=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left[(R+Q-{\widetilde{R}}-{\widetilde{Q}})\partial_{t}{\bf v}+\left((R+Q){\bf u}-({\widetilde{R}}+{\widetilde{Q}}){\bf v}\right)\cdot\nabla{\bf v}\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\Big({p_{+}}({\widetilde{\varrho}_{+}})-{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\widetilde{\varrho}_{+}}-{\varrho_{+}})-{p_{+}}({\varrho_{+}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)\Big({p_{-}}({\widetilde{\varrho}_{-}})-{p_{-}}^{\prime}({\widetilde{\varrho}_{-}})({\widetilde{\varrho}_{-}}-{\varrho_{-}})-{p_{-}}({\varrho_{-}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})(\alpha{\varrho_{+}}-\beta{\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{-}}^{\prime}({\widetilde{\varrho}_{-}})\Big((1-\alpha){\varrho_{-}}-(1-\beta){\widetilde{\varrho}_{-}}\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\partial_{t}\alpha+{\bf v}\cdot\nabla\alpha\right){p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}(1-\alpha)+{\bf v}\cdot\nabla(1-\alpha)\Big){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta){p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t\\ &+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}(1-\beta)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\frac{\gamma-1}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\\ &=:J_{1}+\ldots+J_{10}.\end{split}$ (4.1)
5 The weak–strong uniqueness principle
We first notice that
$\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})(\tau)$
$\qquad\geq C\begin{cases}&\displaystyle\int_{\Omega}\Big[\alpha\left({\varrho_{+}}-{\widetilde{\varrho}_{+}}\right)^{2}+(1-\alpha)\left({\varrho_{-}}-{\widetilde{\varrho}_{-}}\right)^{2}\Big]\ {\rm d}{x},\quad{\varrho_{+}},{\varrho_{-}}\in\left[c_{\star},c^{\star}\right],\\[12.91663pt] &\displaystyle\int_{\Omega}\left(1+\alpha{\varrho_{+}}^{\gamma^{+}}+(1-\alpha){\varrho_{-}}^{\gamma^{-}}\right)\ {\rm d}{x},\quad\text{ otherwise},\end{cases}$ (5.1)
for any $c_{\star},c^{\star}>0$ and $\tau\in(0,T)$. Motivated by this observation, we also define the essential and residual sets corresponding to ${\varrho_{+}},{\varrho_{-}}$ as follows
$\Omega_{\rm ess}(t):=\{x\in\Omega:{\varrho_{+}},{\varrho_{-}}\in\left[c_{\star},c^{\star}\right]\},\quad\Omega_{\rm res}(t)=\Omega\backslash\Omega_{\rm ess}(t).$
Let us now more precisely formulate our main theorem comparing global weak solutions on the time interval $(0,T)$ with the local strong solutions.
Theorem 5.1
Let $(\alpha,\varrho_{-},\varrho_{+},{\bf u})$ be a weak solution emanating from the data $(\alpha_{0},\varrho_{-,0},\varrho_{+,0},{\bf u}_{0})$, given by Theorem 2.2 and let $(\beta,{\widetilde{\varrho}_{-}},{\widetilde{\varrho}_{+}},{\bf v})$ be the strong solution emanating from the data $(\beta_{0},\widetilde{\varrho}_{-,0},\widetilde{\varrho}_{+,0},{\bf v}_{0})$. Then there exists a positive constant $C$ depending on the strong solution (but not on the regularity of the weak solution), such that for a.a. $\tau\in(0,T)$
${\cal E}(\alpha,R,Q,{\bf u}\,|\,\beta,{\widetilde{R}},{\widetilde{Q}},{\bf v})(\tau)+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t$
$\qquad\qquad\qquad\leq C{\cal E}(\alpha,R,Q,{\bf u}\,|\,\beta,{\widetilde{R}},{\widetilde{Q}},{\bf v})(0),$
where ${\cal E}(\alpha,R,Q,{\bf u}\,|\,\beta,{\widetilde{R}},{\widetilde{Q}},{\bf v})$ is given in (LABEL:rel_entr).
In particular, if $(\beta_{0},\widetilde{\varrho}_{-,0},\widetilde{\varrho}_{+,0},{\bf v}_{0})=(\alpha_{0},\varrho_{-,0},\varrho_{+,0},{\bf u}_{0})$, then
$(\alpha,\varrho_{-},\varrho_{+},{\bf u})(\tau,x)=(\beta,,{\widetilde{\varrho}_{-}},{\widetilde{\varrho}_{+}},{\bf v})(\tau,x)$
for a.a. $\tau\in(0,T),\,x\in\Omega$.
Before proving Theorem 5.1, we will state and prove an auxiliary result targeting specifically the difference of the volumetric fractions $\alpha$ and $\beta$.
Lemma 5.2
Let $(\alpha,{\bf u})$, $(\beta,{\bf v})$ satisfying
$\alpha,\beta\in L^{\infty}((0,T)\times\Omega)\cap C([0,T];L^{1}(\Omega)),\quad{\bf u},{\bf v}\in L^{2}(0,T;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}))$
be two distributional solutions to the transport equation (3.4). Suppose, in addition that
$\nabla\beta\in L^{\infty}((0,T)\times\Omega;\mathbb{R}^{3}),\quad\operatorname{div}{\bf v}\in L^{1}(0,T;L^{\infty}(\Omega)).$
Then, for any $\delta>0$ and for all $\tau\in[0,T]$, we have
$\begin{split}\int_{\Omega}(\alpha-\beta)^{2}\ {\rm d}{x}\Big|_{0}^{\tau}\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|^{2}_{W^{1,2}(\Omega)}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\int_{\Omega}(\alpha-\beta)^{2}\ {\rm d}{x}\,{\rm d}t.\end{split}$ (5.2)
Proof.  To justify this lemma rigorously, one may use the DiPerna-Lions regularizing technique [6], see also Lemma 7.3 in [15] for the same argument applied to the pure transport equations for $\alpha,\beta$. In our case the equations of $\alpha,\beta$ contain additional nonlinear terms, so the analysis is more involved. We have
$\partial_{t}(\alpha-\beta)+{\bf v}\cdot\nabla(\alpha-\beta)$
$\qquad\qquad=({\bf v}-{\bf u})\cdot\nabla\alpha-\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\alpha\operatorname{div}{\bf u}+\frac{(\gamma-1)(1-\beta)}{\gamma(1-\beta)+\beta}\beta\operatorname{div}{\bf v}.$
Multiplying this equation by $(\alpha-\beta)$, we obtain
$\int_{\Omega}\frac{1}{2}(\alpha-\beta)^{2}\ {\rm d}{x}-\frac{1}{2}\int_{\Omega}\operatorname{div}{\bf v}(\alpha-\beta)^{2}\ {\rm d}{x}$
$\quad=\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla\alpha(\alpha-\beta)\ {\rm d}{x}$
$\qquad+\int_{\Omega}(\beta-\alpha)(\omega_{\alpha}\operatorname{div}{\bf u}-\omega_{\beta}\operatorname{div}{\bf v})\ {\rm d}{x}$
$\quad=-\frac{1}{2}\int_{\Omega}\operatorname{div}({\bf v}-{\bf u})(\alpha-\beta)^{2}\ {\rm d}{x}$
$\qquad+\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla\beta(\alpha-\beta)\ {\rm d}{x}$
$\qquad+\int_{\Omega}(\beta-\alpha)(\omega_{\alpha}-\omega_{\beta})\operatorname{div}{\bf v}\ {\rm d}{x}$
$\qquad+\int_{\Omega}(\beta-\alpha)\omega_{\alpha}(\operatorname{div}{\bf u}-\operatorname{div}{\bf v})\ {\rm d}{x}.$
A straightforward calculation shows
$|\omega_{\alpha}-\omega_{\beta}|\leq C|\alpha-\beta|,$
where $C=C(\|\alpha\|_{L^{\infty}((0,T)\times\Omega)},\|\beta\|_{L^{\infty}((0,T)\times\Omega)})>0$. Indeed,
$\omega_{\alpha}^{\prime}=\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}(1-\alpha)-\frac{\gamma-1}{\gamma(1-\alpha)+\alpha}\alpha-\alpha(1-\alpha)\frac{\gamma-1}{[\gamma(1-\alpha)+\alpha]^{2}}(-\gamma+1),$
yielding
$|\omega_{\alpha}^{\prime}|\leq 2\frac{\gamma+1}{\min\{\gamma,1\}}+\frac{(\gamma+1)^{2}}{(\min\{\gamma,1\})^{2}}.$
We may choose $\delta>0$ arbitrarily small, such that
$\int_{\Omega}\frac{1}{2}(\alpha-\beta)^{2}\ {\rm d}{x}$
$\quad\leq\delta\|\operatorname{div}({\bf v}-{\bf u})\|^{2}_{L^{2}(\Omega)}+\delta\|{\bf v}-{\bf u}\|^{2}_{L^{2}(\Omega)}$
$\quad+C_{\delta}\left(\|\alpha-\beta\|_{L^{\infty}(\Omega)}^{2}+\|\nabla\beta\|_{L^{\infty}(\Omega)}^{2}+\|\operatorname{div}{\bf v}\|^{2}_{L^{\infty}(\Omega)}\right)\int_{\Omega}(\alpha-\beta)^{2}\ {\rm d}{x},$
where we also used the crucial fact that $\omega_{\alpha}$ is uniformly bounded. The statement thus follows from integrating the above inequality with respect to $t\in(0,\tau)$. $\Box$
With Lemma 5.2 at hand, we may proceed to the proof of Theorem 5.1. We will prove it by estimating $J_{1},\ldots,J_{10}$ term by term.
Estimate of $J_{1}$. We split the first term of the remainder (LABEL:remainder_c) as follows
$J_{1}=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left[(R+Q-{\widetilde{R}}-{\widetilde{Q}})\partial_{t}{\bf v}+\left((R+Q){\bf u}-({\widetilde{R}}+{\widetilde{Q}}){\bf v}\right)\cdot\nabla{\bf v}\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t$
$=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left[(R-{\widetilde{R}}+Q-{\widetilde{Q}})(\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v})\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t$
$+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(R+Q)({\bf u}-{\bf v})\cdot\nabla{\bf v}\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t.$
The second term might be estimated by $\int_{0}^{\tau}C(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t$ with $C$ depending on $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}$. To estimate the first term, we use
$\begin{split}R-{\widetilde{R}}=\alpha({\varrho_{+}}-{\widetilde{\varrho}_{+}})-{\widetilde{\varrho}_{+}}(\beta-\alpha),\end{split}$ (5.3)
and similarly for $Q-{\widetilde{Q}}$. Thus together we obtain
$|R-{\widetilde{R}}+Q-{\widetilde{Q}}|\leq C\Big(\sqrt{\alpha}|{\varrho_{+}}-{\widetilde{\varrho}_{+}}|+\sqrt{1-\alpha}|{\varrho_{-}}-{\widetilde{\varrho}_{-}}|+|\beta-\alpha|\Big),$
where $C>0$ depends only on $\|({\widetilde{\varrho}_{+}},{\widetilde{\varrho}_{-}})\|_{L^{\infty}(\Omega)}$ and on $\|\alpha\|_{L^{\infty}(\Omega)}$. Therefore
$\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\mathbf{1}_{\rm ess}\left[(R-{\widetilde{R}}+Q-{\widetilde{Q}})(\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v})\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t$
$\quad\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{L^{2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\Big(\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t,$
with $C_{\delta}$ depending additionally on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}$. For the residual part of this term in $J_{1}$, considering the small and large values of ${\varrho_{-}},{\varrho_{+}}$ separately, we get
$\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\mathbf{1}_{\rm res}\left[(R-{\widetilde{R}}+Q-{\widetilde{Q}})(\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v})\right]\cdot({\bf v}-{\bf u})\ {\rm d}{x}\,{\rm d}t$
$\quad\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{L^{2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t$
$\qquad+\int_{0}^{\tau}\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}\|R|_{\rm res},Q|_{\rm res}\|_{L^{1}(\Omega)}^{\frac{1}{2}}\|(R+Q)|{\bf u}-{\bf v}|^{2}\|_{L^{1}(\Omega)}^{\frac{1}{2}}\,{\rm d}t$
$\quad\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{L^{2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t,$
with $C_{\delta}$ proportional to $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}^{2}$. Ultimately we have
$J_{1}\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{L^{2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t,$
where $C_{\delta}$ depends on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}$ and on $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}$.
Estimate of $J_{2}$ and $J_{3}$. Again, both the two terms will be considered on the essential and residual parts of the domain. Using Taylor expansion for $J_{2}$, we obtain
$\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\mathbf{1}_{\rm ess}\alpha\Big({p_{+}}({\widetilde{\varrho}_{+}})-{p_{+}}^{\prime}({\widetilde{\varrho}_{+}})({\widetilde{\varrho}_{+}}-{\varrho_{+}})-{p_{+}}({\varrho_{+}})\Big)\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$\leq C\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\mathbf{1}_{\rm ess}\alpha|{\widetilde{\varrho}_{+}}-{\varrho_{+}}|^{2}\ {\rm d}{x}\,{\rm d}t\leq\int_{0}^{\tau}C(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t,$
where $C$ depends on $\|\alpha\|_{L^{\infty}(\Omega)}$ and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}$. The estimate for the essential part of $J_{3}$ is similar. The residual part is bounded straight from (5). Summarizing, we have
$J_{2}+J_{3}\leq\int_{0}^{\tau}C(t)\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})\,{\rm d}t.$
Estimate of $J_{4}$ and $J_{5}$. Both terms are again treated in an analogous way, so we estimate only one of them. Using the decomposition (5.3), we easily verify that
$J_{4}$ $=\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\nabla{H_{+}}^{\prime}({\widetilde{\varrho}_{+}})(\alpha{\varrho_{+}}-\beta{\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{L^{2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\Big(\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t,$
where $C_{\delta}$ depends on $\|{\widetilde{\varrho}_{+}}\|_{W^{1,\infty}(\Omega)}$. Moreover,
$J_{5}\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{L^{2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\Big(\bar{\cal E}(R,Q,{\bf u}|{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t,$
with $C_{\delta}$ depending on $\|{\widetilde{\varrho}_{-}}\|_{W^{1,\infty}(\Omega)}$.
Estimate of $J_{6}$, $J_{7}$ and $J_{8}$. Here, all the terms vanish thanks to the algebraic closure. Indeed, we have
$J_{6}+J_{7}=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha\right){p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}(1-\alpha)+{\bf u}\cdot\nabla(1-\alpha)\Big){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\left(\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha\right){p_{+}}({\widetilde{\varrho}_{+}})\ {\rm d}{x}\,{\rm d}t$
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\Big(\partial_{t}\alpha+{\bf u}\cdot\nabla\alpha\Big){p_{-}}({\widetilde{\varrho}_{-}})\ {\rm d}{x}\,{\rm d}t$
$=$  $\,0,$
[…]
$J_{8}=$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta\,{p_{+}}({\widetilde{\varrho}_{+}})+\nabla(1-\beta)\,{p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t$
$=$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}({\bf v}-{\bf u})\cdot\Big(\nabla\beta\,{p_{+}}({\widetilde{\varrho}_{+}})-\nabla\beta\,{p_{-}}({\widetilde{\varrho}_{-}})\Big)\ {\rm d}{x}\,{\rm d}t$
$=$  $\,0.$
Estimate of $J_{9}$ and $J_{10}$. To estimate the remaining terms, we couple them with terms
$\begin{split}-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t\end{split}$
and
$\begin{split}\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{-}}({\widetilde{\varrho}_{-}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t,\end{split}$
[…]
At first, we transform
$J_{9}$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\frac{(\gamma-1)(1-\beta)}{\gamma(1-\beta)+\beta}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{+}}({\widetilde{\varrho}_{+}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$=$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\left[{p_{+}}({\widetilde{\varrho}_{+}})-({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\right]\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\left(\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}-\frac{(\gamma-1)(1-\beta)}{\gamma(1-\beta)+\beta}\right)({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$=:$ $\,L_{1}+L_{2}.$
Notice that the last term $L_{2}$ can be already estimated, because the difference in the bracket is controlled by $C|\alpha-\beta|$, with $C>0$ depending on $\|(\alpha,\beta)\|_{L^{\infty}(\Omega)}$. We thus have
$L_{2}\leq\int_{0}^{\tau}C(t)\Big(\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t,$
with $C>0$ depending additionally on $\|{\widetilde{\varrho}_{+}}\|_{L^{\infty}(\Omega)}$, and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}$.
We now transfer similarly
$J_{10}$ $+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{-}}({\widetilde{\varrho}_{-}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$=$ $-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\frac{(\gamma-1)\beta}{\gamma(1-\beta)+\beta}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha{p_{-}}({\widetilde{\varrho}_{-}})\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$=$ $\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}[{p_{-}}({\widetilde{\varrho}_{-}})-({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})]\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}(1-\alpha)\left(\frac{(\gamma-1)\alpha}{\gamma(1-\alpha)+\alpha}-\frac{(\gamma-1)\beta}{\gamma(1-\beta)+\beta}\right)({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})\operatorname{div}{\bf v}\ {\rm d}{x}\,{\rm d}t$
$=:$ $\,{\emph{\L }}_{1}+{\emph{\L }}_{2}.$
The term ${\emph{\L }}_{2}$ can be estimated by the same way as $L_{2}$, leading to
${\emph{\L }}_{2}\leq\int_{0}^{\tau}C(t)\Big(\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t,$
with $C$ depending on $\|{\widetilde{\varrho}_{-}}\|_{L^{\infty}(\Omega)}$, and $\|\operatorname{div}{\bf v}\|_{L^{\infty}(\Omega)}$.
Finally let us look at the terms $L_{1}$ and ${\emph{\L }}_{1}$. Using again the equality of pressures, we can write
$L_{1}+{\emph{\L }}_{1}$
$\quad=-\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\Big({p_{+}}({\widetilde{\varrho}_{+}})-({\widetilde{\varrho}_{+}}-{\varrho_{+}}){p_{+}}^{\prime}({\widetilde{\varrho}_{+}})-{p_{+}}({\varrho_{+}})\Big)\ {\rm d}{x}\,{\rm d}t$
$\qquad\,\,+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}\alpha\frac{(\gamma-1)(1-\alpha)}{\gamma(1-\alpha)+\alpha}\operatorname{div}{\bf v}\Big({p_{-}}({\widetilde{\varrho}_{-}})-({\widetilde{\varrho}_{-}}-{\varrho_{-}}){p_{-}}^{\prime}({\widetilde{\varrho}_{-}})-{p_{-}}({\varrho_{-}})\Big)\ {\rm d}{x}\,{\rm d}t,$
and these two integrals can be treated exactly as the terms $J_{2},\ J_{3}$ before. Summarizing, we have shown that
$J_{9}+J_{10}=$ $\,L_{1}+L_{2}+{\emph{\L }}_{1}+{\emph{\L }}_{2}$
$\leq$ $\,\int_{0}^{\tau}C(t)\Big(\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t.$
[…]
$\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})\Big|_{0}^{\tau}+\int_{0}^{\tau}\!\!\!\!\int_{\Omega}{\bf S}(\nabla{\bf u}-\nabla{\bf v}):\nabla({\bf u}-{\bf v})\ {\rm d}{x}\,{\rm d}t$
$\quad\leq\delta\int_{0}^{\tau}\|{\bf v}-{\bf u}\|_{W^{1,2}(\Omega)}^{2}\,{\rm d}t+\int_{0}^{\tau}C_{\delta}(t)\Big(\bar{\cal E}(R,Q,{\bf u}\,|\,{\widetilde{R}},{\widetilde{Q}},{\bf v})+\|\alpha-\beta\|_{L^{2}(\Omega)}^{2}\Big)\,{\rm d}t,$
for $C_{\delta}$ depending on $\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{\infty}(\Omega)}$, $\|\nabla{\bf v}\|_{L^{\infty}(\Omega)}$, $\|{\widetilde{\varrho}_{+}}\|_{W^{1,\infty}(\Omega)}$ and $\|{\widetilde{\varrho}_{-}}\|_{W^{1,\infty}(\Omega)}$, as well as on the upper bounds of $\alpha$ and $\beta$, which are equal to $1$.
Summing this estimate up with (5.2) and using the Grönwall argument, we finish the proof of Theorem 5.1. In particular, our main Theorem 2.1 follows. $\Box$
Acknowledgement. The work of Y.L. was supported by National Natural Science Foundation of China (12571228), Natural Science Foundation of Anhui Province (2408085MA018). The work of E.Z. was supported by the EPSRC Early Career Fellowship no. EP/V000586/1.
Data Availability. Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Conflicts of interest. All authors certify that there are no conflicts of interest for this work.
Publishing licence. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.